X company is starting a new merchandising business and provides the following budgets for its two products:

product Revenue total CM

A $349,716 $111554

B $258,973 69,498

New year's budgeted fixed costs are $210,000. X company would like to at least break even in its first year of operations, what must total sales be in order for that to happen [round unit numbers to two decimal places}? Assume that the budgeted product mix will not change

A $297538

B $395726

C $526316

D $700000

E $931000

F $1238230
Answer #1

Total Contribution margin = 111554+69498"= 181052

Total revenue = 349716+258973 = 608689

Contribution margin ratio = 181052/608689 = 30% (rounded)

Breakeven sales = Fixed cost/Contribution margin ratio

= 210,000/30%

= 700,000